The speed of train A is 10 mph slower than train B. Train A travels 210 miles in the same time train B travels 260. What is the speeds of trains A & B ?
Answer by mananth(16946)

train B xmph
train A = x-10 mph
..
train A 210 miles
train B 260 miles
..
Time taken by both the trains is same
..
260/x =210/x-10
260(x-10)=210x
260x-2600=210x
50x=2600
x=52 mph speed of train B
train A speed = 42mph